# Hamiltonian and Liouvillian learning in weakly-dissipative quantum many-body systems

### v1

@@ II.3 Learning the overall scale, and conserved quantities
When imposing Eq. (4) on numerous initial states in quench dynamics, it leads to a large number of constraints on the Hamiltonian and Liouvillian. However, it is worth noting that these conditions are not enough to uniquely specify the Hamiltonian. The conditions in Eq. (4) define the Hamiltonian only up to a scalar factor since they are linear in $H$. Additionally, these conditions apply to any operator $Q$ such that $[Q,H]=0$, which means they define a linear subspace within the space of Hermitian operators that contains all conserved quantities of $\mathcal{L}_{H}$. An ansatz, $A(\bm{c})$, then may contain the Hamiltonian, $H$, and other conserved quantities of $\mathcal{L}_{H}$ which admit a decomposition in the form of $A(\bm{c})$. For instance, if we choose our ansatz to be $k$-local, it contains at most $k$-local conserved quantities. An example, which we will also consider later, is the total magnetization in ion-trap experiments, $Q\sim\sum_{k}\sigma^{z}_{k}$, which is a sum of local operators.
@@
So let us assume that an ansatz contains two linearly independent conserved quantities $H$, and $Q$, corresponding to two linearly independent vectors $\bm{c}^{H}$, and $\bm{c}^{Q}$, the generalization to more conserved quantities is straightforward. Then, by singular value decomposition, one obtains two right singular vectors $\{\bm{v},\bm{v}_{\perp}\}$, which both satisfy Eq. (7). Therefore, the corresponding singular values, $\lambda_{1}$ and $\lambda_{2}$, are degenerate, i.e., $\lambda_{1}=\lambda_{2}$. However, the right singular vectors must not necessarily correspond to $\bm{c}^{H}$ and $\bm{c}^{H}_{\perp}$. Thus, in the presence of conserved quantities, naively solving Eq. (9) would in general reconstruct the parameters of a linear combination of conserved quantities.
@@
To single out $\bm{c}^{H}$ from this subspace, one needs to impose additional constraints that can only be fulfilled by $H$, but not by other conserved quantities, including scalar multiples of $H$. Such a constraint can be obtained from Ehrenfest’s theorem in Eq. (3) by inserting the ansatz $A(\bm{c})$ for $H$, and considering a generic observable, $O$, not commuting with $H$. This leads to additional constraints, which can only be satisfied by the Hamiltonian, and thus $\bm{c}^{H}$. We spell out the modified equations in Appendix A. The optimization problem in Eq. (9) is then modified to
@@
where $M^{\rm add}$, and $\bm{b}(\bm{d})$ contain additional constraints determining the overall scale, and to single out $\bm{c}^{H}$ (see also Appendix A for explicit formulae for these constraints). Both can be estimated from quench experiments described above, requiring only measurements of few-body Pauli operators. Moreover, the auxiliary parameter $\xi$ controls the penalty that is added for violating the condition determining $H$ within the subspace of conserved quantities of $\mathcal{L}_{H}$. Choosing this parameter large enough allows us to reconstruct a $\bm{c}^{\rm rec}\approx\bm{c}^{H}$, with the correct overall scale, as a unique solution of Eq. (11). We emphasize, that by choosing the appropriate additional constraints, Eq. (11) can also be used to learn other conserved quantities of $\mathcal{L}_{H}$.
II.4 Learning the full Liouvillian
@@
Since we have dissipative dynamics, one expects that accurately learning the Hamiltonian also requires learning all the individual dissipation rates. This is, however, not necessarily the case. In some cases, different dissipative processes give the same contribution to Eq. (4), and therefore, cannot be distinguished by the conditions in Eq. (4). This is a unique feature of our approach as it enables us to learn the Hamiltonian $H$, without having to learn the complete Liouvillian $\mathcal{L}_{D}$.
This feature can be understood already with a simple example of a single spin with Hamiltonian $H=\sigma_{y}$, and Liouvillian
[…]
For simplicity, inserting the above Hamiltonian and Liouvillian as an ansatz in the condition in Eq. (4), leads to the following;
$\expectationvalue{\sigma_{y}}_{T}-\expectationvalue{\sigma_{y}}_{0}=\frac{1}{2}\quantity[\sum_{k}\gamma_{k}]\int_{0}^{T}-4\expectationvalue{\sigma_{y}}_{t}\,\mathrm{d}t,$ (13)
where the right-hand side only depends on the sum of dissipation rates. Moreover, the same condition is obtained for a Liouvillian with only $\sigma_{x}$, or $\sigma_{z}$ Lindblad operators. This has two consequences. Firstly, using these conditions one can learn the Hamiltonian without resolving individual dissipation rates. And secondly, one can learn the Hamiltonian with an insufficient ansatz for the Liouvillian. While this is only a simple illustrative example, we note that this also happens for experimentally relevant systems, such as, for instance, the long-range XY model, as we will illustrate in Section III.2.
@@
More generally, the above feature can be summarized as follows. When inserting an ansatz into the conditions in Eq. (4), with the property discussed above, the resulting constraint matrices $\{M^{(k)}\}$ in Eq. (8), where $k$ labels the different Lindblad operators, become linearly dependent. Then, the decomposition $M_{D}(\bm{d})=\sum_{k}d_{k}M^{(k)}$ is not unique. This dependence leads to symmetries in the cost function in Eq. (9), such that the Hamiltonian $H$ can be learned exactly, for different dissipation rates $\bm{d}^{\mathrm{rec}}$. In order to learn the full Liouvillian and resolve individual dissipation rates, one needs to add additional constraints in the same way as one does for excluding conserved quantities and learning overall scale in Section II.3. This can then be phrased as an optimization problem of the form of Eq. (11), which we explain in detail in Appendix A. Moreover, we will show how to choose additional constraints to learn the full Liouvillian in Section III.2.
II.5 Effect of shot noise
In case the ansatz for the dynamics is sufficient, the only source of error is shot noise. Therefore, with a limited measurement budget one obtains a noisy estimate $\widetilde{M}(\bm{d})=M(\bm{d})+E(\bm{d})$, with an error matrix $E(\bm{d})$. Weyl’s inequality states the following bound on the perturbed singular value $\lambda_{1}$
$\lambda_{1}[\widetilde{M}(\bm{d})]\leq\norm{E(\bm{d})}.$ (14)
Therefore, in early stages of the learning procedure, one expects $\lambda_{1}$ to decrease $\sim N_{\rm runs}^{-1/2}$ as we increase the number of runs, $N_{\rm runs}$, in the experiment. Moreover, if an ansatz is insufficient, $\lambda_{1}$ is strictly bounded away from zero, even in the absence of shot noise. Thus, for $N_{\rm runs}$ sufficiently large, $\lambda_{1}$ will reach a plateau. Here, the dominant source of error will be systematic errors due to missing terms in our ansatz.
$|\sin(\theta)|\leq\frac{\norm{E(\bm{d})}}{\delta}.$ (15)
Here, $\delta=\lambda_{2}-\lambda_{1}^{\rm exact}$, and $\lambda_{1}^{\rm exact}=0$ for a sufficient ansatz. In case the elements of $E$ are i.i.d. Gaussian random variables one can establish average bounds on $\sin(\theta)$ [35]. However, we emphasize that the noise will in general be correlated due to the fact that we perform many of the required measurements in parallel. Moreover, the error matrix $E(\bm{d})$ is experimentally inaccessible. Nevertheless, we will demonstrate in the following, that the ratio $\lambda_{1}/\lambda_{2}$, which can be directly computed from experimental data, serves as a quantity to asses an ansatz. This figure of merit is motivated by the fact that $\lambda_{1}$ quantifies the violation of generalized energy conservation in Eq. (4) and that a larger gap $\delta=\lambda_{2}$ improves the upper bound in Eq. (15). As we will see below, the gap is typically larger for ansätze with fewer parameters.
II.6 Strategies for Hamiltonian and Liouvillian learning
@@
A fundamental tension in Hamiltonian and Liouvillian learning is the tradeoff between the number of parameters of an ansatz, and the number of measurements that are required to estimate these parameters from experimental data. While complex ansätze, comprising many parameters, allows us to learn many aspects of the dynamics, they require a large measurement budget to reduce shot noise. In contrast, in the case of a limited measurement budget, it is quite useful to reduce the number of parameters to just a few relevant ones. This is also physically justified as many experimentally relevant Hamiltonians can be effectively characterized by only a few parameters, which may not even increase with system size. Illustrative examples include translationally invariant systems, or systems exhibiting algebraically decaying interactions. While ansätze with few parameters only require a small measurement budget, they are insufficient in the limit of an infinite number of measurements. However, as long as these insufficiencies are small, learning will be limited by shot noise. Therefore, our goal in the following is to find an ansatz comprising only few parameters and the dominant terms of the Hamiltonian and Liouvillian, which do not show any insufficiencies below our limited measurement budget.
To find this ansatz, we compare different ansätze by varying their operator content as well as their parametrization (see also Fig. 1(4) for an illustration of the following discussion). We emphasize that reparametrization can be done solely by classical post-processing of the data. To compare different ansätze, we will use the ratio $\lambda_{1}/\lambda_{2}$ as a quantifier for the learning error of an ansatz. One expects that in the early stages of the learning procedure, this ratio decreases with $N_{\rm runs}^{-1/2}$, as shot noise is the dominant source of error. In this regime, we will consider the reconstructed parameters $\bm{c}^{\rm rec}$ and $\bm{d}^{\rm rec}$, for which we compute error bars using bootstrapping methods. This allows us to identify dominant terms in the Hamiltonian and Liouvillian, thus learning the dominant operator structure. In later stages of the learning procedure, one expects that $\lambda_{1}/\lambda_{2}$ reaches a plateau due to the insufficient of the ansatz. In that case, we extend our ansatz, for instance, by adding additional operators to the Hamiltonian, or additional Lindblad operators to the Liouvillian. When comparing different ansätze, we choose the one which minimizes $\lambda_{1}/\lambda_{2}$. This results in an ansatz containing only few parameters and small missing terms, i.e., with an early shot-noise scaling, and a low plateau.
Reparametrizing an ansatz involves adjusting the structure of its parameters, while leaving the operator content unchanged. For instance, we will use reparametrization to reduce the number of parameters, resulting in a transformation of the vector of Hamiltonian parameters $\bm{c}$ with $n$ entries, to a vector $\bm{c}_{G}$ with with fewer entries, $\tilde{n}\leq n$. A simple way to achieve this is by transforming the vector $\bm{c}$ and the constraint matrices $M_{H}$ and $M_{D}(\bm{d})$ via a parametrization matrix, $G$, encoding dependencies between the parameters. This results in a transformation
$\bm{c}$ $\mapsto\bm{c}_{G}=G^{T}\bm{c},$ (16)
$M_{H}$ $\mapsto M^{H}_{G}=M_{H}G,$ (17)
$M_{D}(\bm{d})$ $\mapsto M^{D}_{G}(\bm{d})=M_{D}(\bm{d})G.$ (18)
Here, $G$ is an isometry from $\mathbb{R}^{\tilde{n}}$ to $\mathbb{R}^{n}$, and $\tilde{n}\leq n$, as $G^{T}G=\mathds{1}_{\tilde{n}\times\tilde{n}}$, and $GG^{T}$ a projector onto the image of $G$ (see Appendix B for details). The projector $GG^{T}$ projects onto a subspace spanned by vectors $\bm{c}$, with $n$ entries, which admit the dependencies encoded in $G$, for instance, translation invariance. Reconstructing the parameters of the reparametrized ansatz then amounts to solving
where the solution $\bm{c}^{\mathrm{rec}}_{G}$ exactly fulfills the dependencies imposed by $G$. Note, that $G$ may also depend (non-linearly) on multiple parameters, and we will consider an example for this in Section III.2.
In certain cases $GG^{T}\bm{c}^{H}\neq\bm{c}^{H}$, i.e., the Hamiltonian does not admit a parametrized form imposed by $G$. In such a case one needs to reconstruct vectors $\bm{c}$ that, although containing $n$ parameters, are nevertheless close to the parametrized ansatz with only $\tilde{n}$ parameters. In such a case, one solves the optimization problem
[…]
where the last term acts as a penalty, or regularization, and its strength is controlled by the regularization parameter $\beta$. Regularization techniques are well established tools in statistics and machine learning, usually employed to prevent overfitting [39]. For $\beta=0$ this reduces to Eq. (9), i.e., $\bm{c}^{\mathrm{rec}}(0)=\bm{c}^{\mathrm{rec}}$. For $\beta>0$ the solution $\bm{c}^{\mathrm{rec}}$ will, to some extend, adhere to the structure imposed by $G$. In the limit $\beta\rightarrow\infty$ one will obtain $\bm{c}^{\mathrm{rec}}(\infty)=\bm{c}^{\mathrm{rec}}_{G}$, see Appendix B. In this case, the regularization acts like the projector $GG^{T}$ onto the image of $G$, where the dependencies encoded in $G$ are exactly fulfilled.
Finally, one notices that given that $GG^{T}\bm{c}_{H}=\bm{c}_{H}$, i.e., the parametrization does not render an ansatz insufficient, the bound in Eq. (15) can only tighten. To see this, one observes that $\lambda_{2}[M_{G}(\bm{d})]\geq\lambda_{2}[M(\bm{d})]$, and $\norm{E(\bm{d})G}\leq\norm{E(\bm{d})}\norm{G}=\norm{E(\bm{d})}$ for any isometry $G$. Therefore, reducing the number of parameters in general increases the gap $\delta=\lambda_{2}$ of the constraint matrix $M(\bm{d})$.
III Numerical Illustrations of Learning Strategies
In this section we illustrate our strategies for Hamiltonian and Liouvillian learning. We will do this in the context of two experimentally relevant spin-Hamiltonians, i.e., the Ising model in Eq. (II.1), and a long-range Hamiltonian with algebraically decaying interactions, see Eq. (28), in the presence of weak dissipation. In the first example, we focus on learning the Hamiltonian, by successively reparametrizing the ansatz based on monitoring $\lambda_{1}/\lambda_{2}$. In the second example, we address learning in the presence of conserved quantities of the Hamiltonian, and collective dissipation.
III.1 Hamiltonian learning in the presence of weak dissipation

### v2

@@ II.3.1 Learning the overall scale, and conserved quantities
The generalized energy conservation conditions in Eq. (4b) define the Hamiltonian only up to a scalar factor since they are linear in $H$. Additionally, these conditions hold for any operator $Q$ with $[Q,H]=0$, i.e., for all conserved quantities of $H$. An ansatz, $A(\bm{c})$, then may contain the Hamiltonian, $H$, and other conserved quantities of $H$ which admit a decomposition in the form of $A(\bm{c})$. For instance, if we choose our ansatz to be $k$-local, it contains at most $k$-local conserved quantities, such as the total magnetization, $Q\sim\sum_{k}\sigma^{z}_{k}$, which is a sum of local operators.
@@
So let us assume that an ansatz contains two linearly independent conserved quantities $H$, and $Q$, corresponding to two linearly independent vectors $\bm{c}^{H}$, and $\bm{c}^{Q}$ (the generalization is straightforward). Then, by singular value decomposition, one obtains two right singular vectors $\{\bm{v},\bm{v}_{\perp}\}$, which both satisfy Eq. (11), and degenerate singular values $\lambda_{1}=\lambda_{2}=0$. However, the right singular vectors must not necessarily correspond to $\bm{c}^{H}$ and $\bm{c}^{H}_{\perp}$. Thus, in the presence of conserved quantities, naively solving Eq. (14) would in general reconstruct the parameters of a linear combination of conserved quantities.
@@
To distinguish $\bm{c}^{H}$ from other conserved quantities, we impose additional constraints that can only be fulfilled by $H$, but not by other conserved quantities, including scalar multiples of $H$. Such a constraint can be obtained from Eq. (4a) for a generic observable, $O$, not commuting with $H$. We spell out the modified equations in Appendix A. The optimization problem in Eq. (14) is then modified to
@@
where $M(\bm{d})=M_{H}+M_{D}(\bm{d})$ is as in Eq. (14), and $M^{\rm add}$ and $\bm{b}(\bm{d})$ contain additional constraints. The auxiliary parameter $\xi$ controls a penalty that is added for violating the condition determining $H$. Choosing this parameter large enough allows us to reconstruct a $\bm{c}^{\rm rec}\approx\bm{c}^{H}$, with the correct overall scale, as a unique solution of Eq. (15). We emphasize, that by choosing the appropriate additional constraints, Eq. (15) can also be used to learn other conserved quantities of $H$.
II.3.2 Learning the full Liouvillian
@@
Since we have dissipative dynamics, one expects that accurately learning the Hamiltonian also requires learning all the individual dissipation rates. This is, however, not necessarily the case. In some cases, different dissipative processes give the same contribution to Eq. (4b), and therefore, cannot be distinguished by these conditions. More specifically, when inserting an ansatz into this condition, the resulting constraint matrices $\{M^{(k)}\}$ in Eq. (13), where $k$ labels the different Lindblad operators, become linearly dependent. Then, the decomposition $M_{D}(\bm{d})=\sum_{k}d_{k}M^{(k)}$ is not unique. This dependence leads to symmetries in the cost function in Eq. (14), such that the Hamiltonian $H$ can be learned exactly, for different dissipation rates $\bm{d}^{\mathrm{rec}}$.
@@
In order to learn the full Liouvillian and resolve individual dissipation rates, one needs to add additional constraints in the same way as one does for excluding conserved quantities and for learning the overall scale. This can then be phrased as an optimization problem of the form of Eq. (15), which we explain in detail in Appendix A. This is a unique feature of Eq. (11). Below, we will compare Hamiltonian and Liouvillian learning based on Eq. (7), which, generically, does not admit such symmetries, and Eq. (11), including additional constraints, where we will show how to choose these additional constraints to learn the full Liouvillian.
III Numerical Case Studies
In this section we want to illustrate the Hamiltonian and Liouvillian learning methods developed in the previous sections in the context of several experimentally relevant spin-Hamiltonians, more specifically the Ising model in Eq. (II.1), and a long-range Hamiltonian with algebraically decaying interactions that is naturally found in trapped ion experiments [9]. We will study the procedure of learning these Hamiltonians, in the presence of weak dissipation, using only a finite amount of numerically simulated measurement data. Moreover, we will aim to learn the Liouvillian, and in particular distinguish local from collective dissipation. As a preparation for the following discussion let us discuss some details of the learning procedure.
III.1 Preliminaries
To learn the Hamiltonian and Liouvillian of a many-body system, the constraint matrices in Eq. (7), or Eq. (11), are estimated from a finite amount of experimental data. Therefore, we have to solve the corresponding optimization problem using their noisy estimates, e.g., $\tilde{K}_{H}=K_{H}+E$, where $E$ is an error matrix with random entries. Crucially, $E$ is experimentally inaccessible. Therefore, as an experimentally accessible figure of merit to asses the progress of our learning procedure, one may consider the minimum of the objective function in Eq. (10), or Eq. (14), given by the smallest singular value $\lambda_{1}$. In early stages of the learning, where shot-noise, represented by the matrix $E$, is the dominant source of error, one expects this minimum to decrease $\sim N_{\rm runs}^{-1/2}$ as we increase $N_{\rm runs}$, see Appendix B for details. In this regime, we will consider the reconstructed parameters $\bm{c}^{\rm rec}$ and $\bm{d}^{\rm rec}$, for which we compute error bars using bootstrapping methods. This allows us to identify dominant terms in the Hamiltonian and Liouvillian, thus learning the dominant operator structure. For a sufficient ansatz, this error will asymptotically converge to zero, while for an insufficient ansatz the error will be strictly bounded away from zero, i.e., it will plateau in later stages of the learning, where missing ansatz terms are the dominant source of error. Therefore, when observing a plateau we need to extend our ansatz by additional terms.
@@ III.1 Preliminaries
A fundamental tension in Hamiltonian and Liouvillian learning is the tradeoff between the number of parameters of an ansatz, and the number of measurements that are required to estimate these parameters from experimental data. While ansätze comprising many parameters allow us to learn many aspects of the dynamics, they require a large measurement budget to reduce shot noise. In contrast, in the case of a limited measurement budget, it is quite useful to reduce the number of parameters to just a few relevant ones. This is also physically justified as many experimentally relevant Hamiltonians can be effectively characterized by only a few parameters, which may not even increase with system size. Illustrative examples include translationally invariant systems, or systems exhibiting algebraically decaying interactions. While ansätze with few parameters only require a small measurement budget, they might be insufficient in the limit of an infinite number of measurements. However, as long as these insufficiencies are small, learning will be limited by shot noise. Therefore, our goal in the following is to find an ansatz comprising only few parameters and the dominant terms of the Hamiltonian and Liouvillian, which do not show any insufficiencies given our limited measurement budget.
To find this ansatz, we compare different ansätze by varying their operator content as well as their parametrization, i.e., the dependencies between parameters. We emphasize that reparametrization can be done solely by classical post-processing of the data. More specifically, by transforming the vector of parameters, $\bm{c}$, via a parametrization matrix $G$ to a new vector of parameters, $\bm{c}_{G}=G^{T}\bm{c}$, where $G$ encodes dependencies between parameters in $\bm{c}$. This leads to a transformation of the learning equations in Eqs. (7), and (11) [see Appendix C for technical details]. We will demonstrate in the following numerical simulations that the ratio of the smallest two singular values of the constraint matrices, $\lambda_{1}/\lambda_{2}$, serves as a quantifier for the learning error of an ansatz and a useful figure of merit to compare different ansätze, . This is motivated by the fact that the smallest singular value, $\lambda_{1}$, quantifies the learning error, e.g., the violation of generalized energy conservation in Eq. (4b), and that a larger gap $\delta=\lambda_{2}$ tightens the upper bound on $|\sin(\theta)|$, where $\theta=\angle(\bm{c}^{H},\bm{c}^{\rm rec})$ is the reconstruction error (see Appendix B for details). As we will see below, the gap is typically larger for ansätze with fewer parameters. Therefore, when comparing different ansätze, we choose the one which minimizes $\lambda_{1}/\lambda_{2}$. This results in an ansatz containing only few parameters and small missing terms, i.e., with an early shot-noise scaling, and a low plateau (see Fig. 2 for an illustration).
As a final remark, let us note that in case the operators $O_{i}$, $h_{j}$ and $a_{k}$ are Pauli operators, the expectation values in Eqs. (9) and (13) evaluate to $\alpha\expectationvalue{h_{j}}$, where $\alpha=0$ if $[l_{k},h_{j}]=0$, and $\alpha=-4$ otherwise, and respectively for $O_{i}$. Therefore, in case the operators $\{O_{i}\}$, $\{h_{j}\}$ are few-body operators these expectation values as well as the commutator $[O_{i},h_{j}]$ remain few-body Paulis. Thus, many of the operators commute and can thus be measured jointly.
Estimating the integrals in Eq. (35) requires measuring time-resolved expectation values of the corresponding operators over the duration of the quench, see Fig. 2. Here, we use the composite Simpson’s rule  [^4] It reads $\int_{0}^{T}\expectationvalue*{P}_{t}\mathrm{d}t\approx\frac{1}{3}\Delta t\left[\expectationvalue*{P}_{0}+4\sum_{k=1}^{N_{t}/2}\expectationvalue*{P}_{t_{2k-1}}+2\sum_{k=1}^{N_{t}/2-1}\expectationvalue*{P}_{t_{2k}}+\expectationvalue*{P}_{T}\right]$. which approximates an integral as a series of parabolic segments between $N_{t}$ equally spaced points $t_{m}=m\cdot\Delta t$, with $m\in 0,\dots,N_{t}$ and distance $\Delta t=T/N_{t}$. In the following, we choose a sufficiently small time step $\Delta t$ to ensure that errors arising from discretizing the integral can be disregarded in comparison to shot-noise errors  [^5] Indeed, for fixed time $T$ one obtains $\int_{0}^{T}\expectationvalue*{h}_{t}\mathrm{d}t=I(\Delta t)+K\Delta t^{4}$, where $I(\Delta t)$ is the integral approximation, and $K$ is a constant. When expressed in terms of $N_{t}$ this reads $I(N_{t})+\tilde{K}/N_{t}^{4}$. However, as $I(N_{t})$ can only be estimated from data, one obtains $I(N_{t})=\tilde{I}(N_{t})+\epsilon$, only up to a statistical error $\epsilon$. As $I(N_{t})$ is linear, the variance of $\epsilon$ roughly scales like $1/(N_{t}N_{s})$, where $N_{s}$ is the number of shots per time-point..
Figure 3: Hamiltonian learning in presence of weak dissipation. (a) The learning error $\lambda_{1}/\lambda_{2}$ as a function of the number of runs, $N_{\rm runs}$, of the simulator for different ansätze described in the main text. The colored triangles indicate the asymptotic values for $N_{\rm runs}\to\infty$. The dashed line follows the learning procedure described in the main text. (b) Snapshots of the (normalized) reconstructed Hamiltonian parameters, $c^{\rm rec}_{j}$, for some of the ansätze described in the main text. The capital letters denote groups of operators in the ansatz, and within each group the site indices of the coefficients are sorted increasingly from left to right, e.g., the coefficients in group $XY$ correspond to the operators $\sigma^{x}_{k}\sigma^{y}_{k+1}+\sigma^{y}_{k}\sigma^{x}_{k+1}$ in Eq. (17), and similarly for all other groups of operators. Black crosses indicate the true parameters of the model Hamiltonian as defined below. The point at which these snapshots are taken in terms of $N_{\mathrm{runs}}$ is also indicated by asterisks in the left panel. All error bars are computed via bootstrapping as explained in Appendix D, using $80$ samples. Model parameters: The Hamiltonian parameters in Eq. (16) are chosen as $B_{x}=4/5B_{z}$. Moreover, the coupling strengths are described by $4$-th order polynomials of the form $J_{i,i+1}^{z}=B_{z}\sum_{l=1}^{5}a_{l}\cdot x_{i,i+1}^{l-1}$, and $J_{i,i+2}^{z}=B_{z}\sum_{l=1}^{5}b_{l}\cdot x_{i,i+2}^{l-1}$ where $x_{i,j}=[(i+j)-(N+1)]/N$. We choose the following coefficients: $\bm{a}=(6/5,1/20,1/5,0,-2/5)$ and $\bm{b}=(1/5,1/20,-2/5,0,4/5)$. The dissipation rates are chosen such that $(\gamma_{+},\gamma_{-},\gamma_{z})=(1,1.5,2)\times 10^{-2}B^{z}\ll\|H\|$. To estimate the integrals in Eq. (13), we use $B_{z}\Delta_{t}=1/64$, i.e., $64$ equidistant times, where at each time we spend only $1/64$ of the measurements per basis, compared to the quench times $B_{z}T\in[0,0.5,1]$. Expectation values at $t=0$ are calculated exactly from the initial state.
III.2 Learning from generalized energy conservation